# The complex structure of the Teichmüller space of circle diffeomorphisms in the Zygmund smooth class

### v2

@@ 3. Pre-Schwarzian derivatives
Let $\psi_{f}(z)=\log f^{\prime}(z)$ for a locally univalent holomorphic function $f$. Its derivative $P_{f}(z)=\psi_{f}^{\prime}(z)$ is called the pre-Schwarzian derivative of $f$. Let $B(\mathbb{D})$ be the space of all holomorphic functions $\psi$ on $\mathbb{D}$ such that
@@
We require that $f_{\mu}$ satisfies $f_{\mu}(\infty)=\infty$. Then, $P_{{f_{\mu}}|_{\mathbb{D}}}$ is well defined for every $\mu\in{\rm Bel}(\mathbb{D}^{*})$ by this normalization and $\psi_{{f_{\mu}}|_{\mathbb{D}}}$ belongs to $B(\mathbb{D})$. We may assume that $\psi_{{f_{\mu}}|_{\mathbb{D}}}(0)=0$. By this correspondence $\mu\mapsto\psi_{{f_{\mu}}|_{\mathbb{D}}}$, we have a holomorphic map $\Psi:{\rm Bel}(\mathbb{D}^{*})\to B(\mathbb{D})$ called the pre-Schwarzian derivative map.
Let $\widetilde{\mathcal{T}}=\Psi({\rm Bel}(\mathbb{D}^{*}))$, which consists of those $\psi_{f}$ for a conformal homeomorphism $f$ of $\mathbb{D}$ that is quasiconformally extendable to $\mathbb{D}^{*}$ with $f(\mathbb{D})$ bounded. It is known that this is an open subset of $B(\mathbb{D})$ but there are other uncountably many connected components of the open subset in $B(\mathbb{D})$ consisting of those $\psi_{f}$ with $f(\mathbb{D})$ unbounded. See [16].
@@
which we regard as the complex Banach space by ignoring the difference of complex constant functions. We note that $\psi\in B^{Z}(\mathbb{D})$ if and only if $\psi^{\prime}\in B(\mathbb{D})$. Moreover, we see that $\|\psi\|_{B}\lesssim\|\psi\|_{B^{Z}}$ (see [17, Theorem 5.4]). Hereafter, the symbol $\lesssim$ stands for the inequality modulo a positive constant multiple. The space $B^{Z}(\mathbb{D})$ was defined in [13] by using $\sup_{z\neq w\in\mathbb{D}}|\psi^{\prime}(z)-\psi^{\prime}(w)|/d(z,w)$ for the hyperbolic distance $d$ on $\mathbb{D}$, but they are equivalent (see [17, Theorem 5.5]).
@@
by [13, Theorem 1.1]. This is an open subset of $B^{Z}(\mathbb{D})$ because $\widetilde{\mathcal{T}}$ is open in $B(\mathbb{D})$ and $\|\cdot\|_{B}\lesssim\|\cdot\|_{B^{Z}}$. Moreover, unlike in the case of the universal Teichmüller space $T$, it is proved in [13, Theorem 1.3] that $\widetilde{\mathcal{T}}^{Z}$ is exactly the set of those $\psi_{f}$ for a conformal homeomorphism $f$ of $\mathbb{D}$ that is quasiconformally extendable to $\mathbb{D}^{*}$ with its complex dilatation in ${\rm Bel}^{Z}(\mathbb{D}^{*})$.
@@
Lemma 12.
The map $\Lambda:B^{Z}(\mathbb{D})\to A^{Z}(\mathbb{D})$ defined by $\Lambda(\psi)=\psi^{\prime\prime}-\frac{1}{2}(\psi^{\prime})^{2}$ is holomorphic.
@@
It is obvious that $\Lambda$ is Gâteaux holomorphic. Hence, to prove that $\Lambda$ is holomorphic, it suffices to show that $\Lambda$ is locally bounded. See [1, p.28]. The local boundedness can be verified as follows.
@@
Here, by [17, Theorem 5.4], we see that
@@
We state the main claim in this section, which improves Theorem 11.
Theorem 13.
Both $\Psi:{\rm Bel}^{Z}(\mathbb{D}^{*})\to\widetilde{\mathcal{T}}^{Z}\subset B^{Z}(\mathbb{D})$ and $\Lambda|_{\widetilde{\mathcal{T}}^{Z}}:\widetilde{\mathcal{T}}^{Z}\to\alpha(T^{Z})\subset A^{Z}(\mathbb{D})$ are holomorphic split submersions.
@@
By Theorem 7, $\Phi=\Lambda\circ\Psi$ is a holomorphic split submersion. Combined with Theorem 11 and Lemma 12, this implies the statement. ∎
In fact, $\widetilde{\mathcal{T}}^{Z}$ is a complex analytic disk-bundle over $T^{Z}$.
@@
For the Teichmüller space $T^{\gamma}$ of circle diffeomorphisms of $\gamma$-Hölder continuous derivatives, the fact that the pre-Schwarzian derivative map $\Psi$ is a holomorphic split submersion was proved in [14, Theorem 1.2]. Once we know that $\Psi$ is holomorphic (in fact, it suffices to prove its continuity to see this), the fact that the Schwarzian derivative map $\Phi$ is a holomorphic split submersion shown in [9, Theorem 7.6] implies that so is $\Psi$ by similar arguments as above.

### v3

@@ 3. Pre-Schwarzian derivatives
Let $\psi=\log f^{\prime}$ for a locally univalent holomorphic function $f$ on $\mathbb{D}$. Its derivative $P_{f}=\psi^{\prime}$ is called the pre-Schwarzian derivative of $f$. Let $B(\mathbb{D})$ be the space of all holomorphic functions $\psi$ on $\mathbb{D}$ such that
@@
We require that $f_{\mu}$ satisfies $f_{\mu}(\infty)=\infty$. Then, $P_{{f_{\mu}}|_{\mathbb{D}}}$ is well defined for every $\mu\in{\rm Bel}(\mathbb{D}^{*})$ by this normalization and $\psi=\log\,({f_{\mu}}|_{\mathbb{D}})^{\prime}$ belongs to $B(\mathbb{D})$. We may assume that $\psi(0)=0$. By this correspondence $\mu\mapsto\log\,({f_{\mu}}|_{\mathbb{D}})^{\prime}$, we have a holomorphic map $\Psi:{\rm Bel}(\mathbb{D}^{*})\to B(\mathbb{D})$ called the pre-Schwarzian derivative map.
Let $\widetilde{\mathcal{T}}=\Psi({\rm Bel}(\mathbb{D}^{*}))$, which consists of those $\psi=\log f^{\prime}$ for a conformal homeomorphism $f$ of $\mathbb{D}$ that is quasiconformally extendable to $\mathbb{D}^{*}$ with $f(\mathbb{D})$ bounded. It is known that this is an open subset of $B(\mathbb{D})$ but there are other uncountably many connected components of the open subset in $B(\mathbb{D})$ consisting of those $\psi=\log f^{\prime}$ with $f(\mathbb{D})$ unbounded. See [17].
@@
which we regard as the complex Banach space by ignoring the difference of complex constant functions. We note that $\psi\in B^{Z}(\mathbb{D})$ if and only if $\psi^{\prime}\in B(\mathbb{D})$. Moreover, we see that $\|\psi\|_{B}\lesssim\|\psi\|_{B^{Z}}$ (see [18, Theorem 5.4]). Hereafter, the symbol $\lesssim$ stands for the inequality modulo a positive constant multiple. The space $B^{Z}(\mathbb{D})$ was defined in [14] by using $\sup_{z\neq w\in\mathbb{D}}|\psi^{\prime}(z)-\psi^{\prime}(w)|/d(z,w)$ for the hyperbolic distance $d$ on $\mathbb{D}$, but they are equivalent (see [18, Theorem 5.5]).
@@
by [14, Theorem 1.1]. This is an open subset of $B^{Z}(\mathbb{D})$ because $\widetilde{\mathcal{T}}$ is open in $B(\mathbb{D})$ and $\|\!\!\!~{}\cdot~{}\!\!\|_{B}\lesssim\|\cdot\|_{B^{Z}}$. Moreover, unlike in the case of the universal Teichmüller space $T$, it is proved in [14, Theorem 1.3] that $\widetilde{\mathcal{T}}^{Z}$ is exactly the set of those $\psi=\log f^{\prime}$ for a conformal homeomorphism $f$ of $\mathbb{D}$ that is quasiconformally extendable to $\mathbb{D}^{*}$ with its complex dilatation in ${\rm Bel}^{Z}(\mathbb{D}^{*})$. In other words, for every $f=f_{\mu}|_{\mathbb{D}}$ with $\mu\in{\rm Bel}^{Z}(\mathbb{D}^{*})$, its image $f(\mathbb{D})$ is a bounded domain in $\mathbb{C}$.
@@
Lemma 3.2.
The map $\Lambda:B^{Z}(\mathbb{D})\to A^{Z}(\mathbb{D})$ defined by $\Lambda(\psi)=\psi^{\prime\prime}-(\psi^{\prime})^{2}/2$ is holomorphic.
@@
It is obvious that $\Lambda$ is Gâteaux holomorphic. Hence, to prove that $\Lambda$ is holomorphic, it suffices to show that $\Lambda$ is locally bounded. See [1, p.28] and [2, Theorem 14.9]. The local boundedness can be verified as follows.
@@
Here, by [18, Theorem 5.4], we see that
@@
We can improve Theorem 3.1 immediately in virtue of Theorem 2.4. It also asserts the same property for $\Lambda=\Lambda|_{\widetilde{\mathcal{T}}^{Z}}$.
Theorem 3.3.
Both $\Psi:{\rm Bel}^{Z}(\mathbb{D}^{*})\to\widetilde{\mathcal{T}}^{Z}\subset B^{Z}(\mathbb{D})$ and $\Lambda:\widetilde{\mathcal{T}}^{Z}\to\alpha(T^{Z})\subset A^{Z}(\mathbb{D})$ are holomorphic split submersions.
@@
By Theorem 2.4, $\Phi=\Lambda\circ\Psi$ is a holomorphic split submersion. Combined with Theorem 3.1 and Lemma 3.2, this implies the statement. ∎
@@
For the Teichmüller space $T^{\gamma}$ of circle diffeomorphisms of $\gamma$-Hölder continuous derivatives, the fact that the pre-Schwarzian derivative map $\Psi$ is a holomorphic split submersion was asserted in [15, Theorem 1.2]. Once we know that $\Psi$ is holomorphic (in fact, it suffices to prove its continuity to see this), the fact that the Schwarzian derivative map $\Phi$ is a holomorphic split submersion shown in [10, Theorem 7.6] implies that so is $\Psi$ by similar arguments as above.
We can examine the structure of $\widetilde{\mathcal{T}}^{Z}$ more closely. In fact, $\widetilde{\mathcal{T}}^{Z}$ is a topological disk bundle over $T^{Z}$ though we do not know whether or not this is biholomorphically equivalent to the Bers fiber space over $T^{Z}$. To see this, we use another pre-Schwarzian derivative map given by a different normalization. For any $\zeta\in\overline{\mathbb{D}^{*}}$, we impose the condition $f_{\mu}(\zeta)=\infty$ on the quasiconformal homeomorphism $f_{\mu}$ conformal on $\mathbb{D}$ with complex dilatation $\mu$ on $\mathbb{D}^{*}$. Then, $P_{{f_{\mu}}|_{\mathbb{D}}}$ is well defined for every $\mu\in{\rm Bel}(\mathbb{D}^{*})$ by this normalization and $\psi=\log({f_{\mu}}|_{\mathbb{D}})^{\prime}$ belongs to $B(\mathbb{D})$. We also assume $f_{\mu}(0)=0$ and $(f_{\mu})^{\prime}(0)=1$ hereafter, and denote such a map $f_{\mu}$ by $f_{\mu}^{\zeta}$. It satisfies $\psi(0)=0$ and $\psi^{\prime}(0)=(f_{\mu}^{\zeta})^{\prime\prime}(0)$. By this correspondence $\mu\mapsto\log(f_{\mu}^{\zeta}|_{\mathbb{D}})^{\prime}$, we have a holomorphic split submersion $\Psi_{\zeta}:{\rm Bel}^{Z}(\mathbb{D}^{*})\to\widetilde{\mathcal{T}}^{Z}\subset B^{Z}(\mathbb{D})$ in the same way as before if $\zeta\in\mathbb{D}^{*}$.
Theorem 3.4.
$\widetilde{\mathcal{T}}^{Z}$ is a topological bundle over $T^{Z}\cong\alpha(T^{Z})$ with the projection $\Lambda:\widetilde{\mathcal{T}}^{Z}\to\alpha(T^{Z})$ and the fiber homeomorphic to $\mathbb{D}^{*}$.
Proof.
For every $\varphi_{0}\in\alpha(T^{Z})=\Phi({\rm Bel}^{Z}(\mathbb{D}^{*}))$, we take some $\mu_{0}\in{\rm Bel}^{Z}(\mathbb{D}^{*})$ such that $\Phi(\mu_{0})=\varphi_{0}$. By Theorem 2.4, there is a holomorphic map $s_{\mu_{0}}:V_{\varphi_{0}}\to{\rm Bel}^{Z}(\mathbb{D}^{*})$ defined on some neighborhood $V_{\varphi_{0}}\subset\alpha(T^{Z})$ of $\varphi_{0}$ such that $s_{\mu_{0}}(\varphi_{0})=\mu_{0}$. Then, we define a local trivialization
$L:V_{\varphi_{0}}\times\mathbb{D}^{*}\to\widetilde{\mathcal{T}}^{Z}$
by $L(\varphi,\zeta)=\Psi_{\zeta}(s_{\mu_{0}}(\varphi))$, which satisfies $\Lambda\circ L(\varphi,\zeta)=\varphi$ for every $\varphi\in V_{\varphi_{0}}$. We will prove that $L$ is a homeomorphism onto $\Lambda^{-1}(V_{\varphi_{0}})$. This implies that $\widetilde{\mathcal{T}}^{Z}$ is a disk bundle over $T^{Z}$.
For any $\mu\in{\rm Bel}^{Z}(\mathbb{D}^{*})$ and for any $\zeta,\zeta_{*}\in\mathbb{D}^{*}$, the difference of $f_{\mu}^{\zeta}$ from $f_{\mu}^{\zeta_{*}}$ is given by the post-composition of a normalized Möbius transformation $W_{c}(z)=z/(cz+1)$ with $W_{c}(0)=0$ and $(W_{c})^{\prime}(0)=1$, where $c$ satisfies $f^{\zeta_{*}}_{\mu}(\zeta)=-1/c$. Namely, $f^{\zeta}_{\mu}=W_{c}\circ f^{\zeta_{*}}_{\mu}$. From this, we see that $L$ is injective. For any $\varphi\in V_{\varphi_{0}}$, we define the restriction of $L$ onto the fiber $\Lambda^{-1}(\varphi)$ as
$\ell_{\varphi}=L|_{\{\varphi\}\times\mathbb{D}^{*}}:\mathbb{D}^{*}\to\Lambda^{-1}(\varphi).$
We show that $\ell_{\varphi}$ is surjective onto $\Lambda^{-1}(\varphi)$ as follows. Let $\mu=s_{\mu_{0}}(\varphi)$. Every $\psi\in\Lambda^{-1}(\varphi)$ is represented as $\psi=\log\,(W_{c}\circ f_{\mu}|_{\mathbb{D}})^{\prime}$ for a Möbius transformation $W_{c}$ with $-1/c\in f_{\mu}(\mathbb{D}^{*})$ because $W_{c}\circ f_{\mu}(\mathbb{D})$ is a bounded domain in $\mathbb{C}$. We can find this fact in [14, Theorem 1.3] and also see it by the argument below. Then, for $\zeta\in\mathbb{D}^{*}$ satisfying $f_{\mu}(\zeta)=-1/c$, we have $\ell_{\varphi}(\zeta)=\log\,(f_{\mu}^{\zeta}|_{\mathbb{D}})^{\prime}=\psi$. Hence, $\ell_{\varphi}$ is surjective for every $\varphi\in V_{\varphi_{0}}$, and so is $L$.
A simple computation yields
$P_{f^{\zeta}_{\mu}|_{\mathbb{D}}}(z)-P_{f^{\zeta_{*}}_{\mu}|_{\mathbb{D}}}(z)$ $=\frac{-2(f^{\zeta_{*}}_{\mu})^{\prime}(z)}{{f^{\zeta_{*}}_{\mu}}(z)-f^{\zeta_{*}}_{\mu}(\zeta)},$
$(P_{f^{\zeta}_{\mu}|_{\mathbb{D}}})^{\prime}(z)-(P_{f^{\zeta_{*}}_{\mu}|_{\mathbb{D}}})^{\prime}(z)$ $=\frac{2(f^{\zeta_{*}}_{\mu})^{\prime}(z)^{2}}{(f^{\zeta_{*}}_{\mu}(z)-f^{\zeta_{*}}_{\mu}(\zeta))^{2}}-\frac{2(f^{\zeta_{*}}_{\mu})^{\prime\prime}(z)}{f^{\zeta_{*}}_{\mu}(z)-f^{\zeta_{*}}_{\mu}(\zeta)}$
for $z\in\mathbb{D}$. To see that $\ell_{\varphi}$ is continuous, it suffices to show that
$|P_{f^{\zeta}_{\mu}|_{\mathbb{D}}}(0)-P_{f^{\zeta_{*}}_{\mu}|_{\mathbb{D}}}(0)|\to 0,\quad\sup_{|z|<1}\,(1-|z|^{2})|(P_{f^{\zeta}_{\mu}|_{\mathbb{D}}})^{\prime}(z)-(P_{f^{\zeta_{*}}_{\mu}|_{\mathbb{D}}})^{\prime}(z)|\to 0$
as $\zeta\to\zeta_{*}$. The former limit follows from $f^{\zeta_{*}}_{\mu}(0)=0$, $(f^{\zeta_{*}}_{\mu})^{\prime}(0)=1$, and $f^{\zeta_{*}}_{\mu}(\zeta)\to\infty$ as $\zeta\to\zeta_{*}$. For the latter limit, we note that $|(f^{\zeta_{*}}_{\mu})^{\prime}(z)|$ is bounded and bounded away from 00 by [10, Proposition 6.8]. Moreover,
$\sup_{|z|<1}\,(1-|z|^{2})|(f^{\zeta_{*}}_{\mu})^{\prime\prime}(z)|=\sup_{|z|<1}\,(1-|z|^{2})|P_{f^{\zeta_{*}}_{\mu}}(z)||(f^{\zeta_{*}}_{\mu})^{\prime}(z)|<\infty.$
Thus, the continuity of $\ell_{\varphi}$ can be verified. In fact, $\ell_{\varphi}$ is equi-continuous when $\varphi$ moves locally and so does $\mu$.
Since $\ell_{\varphi}$ is a continuous bijection, the continuity of $\ell_{\varphi}^{-1}$ follows if we prove that $\ell_{\varphi}(\zeta)$ diverges as $\zeta\in\mathbb{D}^{*}$ tends to the boundary $\partial\mathbb{D}^{*}=\mathbb{S}$. We apply the above computation for $\zeta^{*}=\infty$ ($f_{\mu}^{\infty}=f_{\mu}$) and $\zeta\in\mathbb{S}$ to obtain an inequality
$\quad\ (1-|z|^{2})|(P_{f^{\zeta}_{\mu}|_{\mathbb{D}}})^{\prime}(z)-(P_{f_{\mu}|_{\mathbb{D}}})^{\prime}(z)|$
$\geq\frac{2(1-|z|^{2})|(f_{\mu})^{\prime}(z)|^{2}}{|f_{\mu}(z)-f_{\mu}(\zeta)|^{2}}-\frac{2(1-|z|^{2})|(f_{\mu})^{\prime\prime}(z)|}{|f_{\mu}(z)-f_{\mu}(\zeta)|}$
$=\frac{2|(f_{\mu})^{\prime}(z)|}{|f_{\mu}(z)-f_{\mu}(\zeta)|}\left\{\frac{\delta(f_{\mu}(z))}{|f_{\mu}(z)-f_{\mu}(\zeta)|}-(1-|z|^{2})|P_{f_{\mu}|_{\mathbb{D}}}(z)|\right\}.$
Here, $\delta(w)=(1-|z|^{2})|(f_{\mu})^{\prime}(z)|$ for $w=f_{\mu}(z)$ is the inverse of the hyperbolic density of the quasidisk $f_{\mu}(\mathbb{D})$, which is comparable to the distance from $w$ to the boundary $\partial f_{\mu}(\mathbb{D})$. Moreover, as $z\in\mathbb{D}$ tends to $\zeta\in\mathbb{S}$ radially, $\delta(f_{\mu}(z))\asymp|f_{\mu}(z)-f_{\mu}(\zeta)|$ by the geometric property of quasiconformal mappings. Further, $(1-|z|^{2})|P_{f_{\mu}|_{\mathbb{D}}}(z)|\to 0$ as $|z|\to 1$. Therefore, the formulas in the above inequality diverge to $\infty$ as $z$ tends to $\zeta$ radially.
Since $\sup_{|z|<1}(1-|z|^{2})|(P_{f_{\mu}|_{\mathbb{D}}})^{\prime}(z)|<\infty$, we have $\sup_{|z|<1}(1-|z|^{2})|(P_{f^{\zeta}_{\mu}|_{\mathbb{D}}})^{\prime}(z)|=\infty$ for $\zeta\in\mathbb{S}$. This implies that if $\log\,(f^{\zeta}|_{\mathbb{D}})^{\prime}\in B^{Z}(\mathbb{D})$ then $\zeta\in\mathbb{D}^{*}$, which is an alternative proof of the fact that if $\log\,(W_{c}\circ f_{\mu}|_{\mathbb{D}})^{\prime}\in B^{Z}(\mathbb{D})$ then $W_{c}\circ f_{\mu}(\mathbb{D})$ is a bounded domain in $\mathbb{C}$. We can also see that
$\|\ell_{\varphi}(\zeta)\|_{B^{Z}}\geq\sup_{|z|<1}(1-|z|^{2})|(P_{f^{\zeta}_{\mu}|_{\mathbb{D}}})^{\prime}(z)|\to\infty$
as $\zeta\in\mathbb{D}^{*}$ tends to $\mathbb{S}$ because $\ell_{\varphi}$ extends to $\mathbb{S}$ continuously by allowing the value $\infty$.
We have shown that $\ell_{\varphi}=L|_{\{\varphi\}\times\mathbb{D}^{*}}$ is a homeomorphism for every $\varphi\in V_{\varphi_{0}}$. Finally, we check that the bijection $L$ is a homeomorphism. For the continuity of $L$, we use the triangle inequality
$\|L(\varphi,\zeta)-L(\varphi_{*},\zeta_{*})\|_{B^{Z}}\leq\|L(\varphi,\zeta)-L(\varphi,\zeta_{*})\|_{B^{Z}}+\|L(\varphi,\zeta_{*})-L(\varphi_{*},\zeta_{*})\|_{B^{Z}}.$
Then, for the first term in the right side, we apply the equi-continuity of $\ell_{\varphi}$ to show that the norm converges to 00 as $\zeta\to\zeta_{*}$. For the second term, we apply the continuity of $\Psi_{\zeta_{*}}\circ s_{\mu_{0}}$ to show that the norm converges to 00 as $\varphi\to\varphi_{*}$.
For the continuity of $L^{-1}$, we consider any $\psi=L(\varphi,\zeta)$ and $\psi_{*}=L(\varphi_{*},\zeta_{*})$ in $\Lambda^{-1}(V_{\varphi_{0}})$, and choose $\tilde{\psi}=L(\varphi_{*},\zeta)$. Then, we have
$\|L^{-1}(\psi)-L^{-1}(\psi_{*})\|_{A^{Z}\times\mathbb{D}^{*}}$ $\leq\|L^{-1}(\psi)-L^{-1}(\tilde{\psi})\|_{A^{Z}\times\mathbb{D}^{*}}+\|L^{-1}(\tilde{\psi})-L^{-1}(\psi_{*})\|_{A^{Z}\times\mathbb{D}^{*}}$
$=\|(\varphi,\zeta)-(\varphi_{*},\zeta)\|_{A^{Z}\times\mathbb{D}^{*}}+\|(\varphi_{*},\zeta)-(\varphi_{*},\zeta_{*})\|_{A^{Z}\times\mathbb{D}^{*}}$
$=\|\varphi-\varphi_{*}\|_{A^{Z}}+|\zeta-\zeta_{*}|.$
The definition of the norm $\|\cdot\|_{A^{Z}\times\mathbb{D}^{*}}$ on the product space is obvious and $|\zeta-\zeta_{*}|$ should be interpreted appropriately if either $\zeta$ or $\zeta_{*}$ is $\infty$. The convergence $\psi\to\psi_{*}$ implies that $\varphi\to\varphi_{*}$, and hence $\psi-\tilde{\psi}\to 0$ by the continuity of $L$. This also implies that
$\|\ell_{\varphi_{*}}(\zeta)-\ell_{\varphi_{*}}(\zeta_{*})\|_{B^{Z}}=\|\tilde{\psi}-\psi_{*}\|_{B^{Z}}\leq\|\tilde{\psi}-\psi\|_{B^{Z}}+\|\psi-\psi_{*}\|_{B^{Z}}\to 0.$
By the continuity of $\ell_{\varphi_{*}}^{-1}$, we have that $\zeta\to\zeta_{*}$. Thus, we obtain that $L^{-1}$ is continuous, and $L$ is a homeomorphism. ∎
Remark 5.
For the Teichmüller space $T^{\gamma}$ of circle diffeomorphisms of $\gamma$-Hölder continuous derivatives, the same result holds true for the fiber space $\widetilde{\mathcal{T}}^{\gamma}$ defined in the same manner.